OpenStudy (anonymous):

[6.02] Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1

OpenStudy (anonymous):

i will help u if u help me

OpenStudy (dtan5457):

I would definitely isolate y in the first equation.

OpenStudy (dtan5457):

2x+y=3 y=3-2x

OpenStudy (dtan5457):

Plug that into the 2nd equation and tell me what you get.

OpenStudy (anonymous):

could u help me @dtan5457

OpenStudy (anonymous):

-3x + 4(3-2x)=1

OpenStudy (dtan5457):

Good. Solve for x now.

OpenStudy (anonymous):

HELPPPPPPP

OpenStudy (anonymous):

x=1?

OpenStudy (dtan5457):

Yes!

OpenStudy (dtan5457):

Now that you have your x value, you need to find your y value. 2x + y = 3 that x=1 into that.

OpenStudy (dtan5457):

Now solve for y.

OpenStudy (anonymous):

Ok i got it from here thank you!

OpenStudy (dtan5457):

Welcome :)